A volume of 2000 mL is to infuse at 175 mL/hr using an IV set with drop factor of 15 gtt/mL. The infusion was started at 21:35. Calculate : a) gtt/min; b ) infusion time; c) completion time
Answer #1

ANSWER

a) gtt/min.

         Formula = mL /hr * drop factor / 60 minutes

                Volume = 2000ml.

              Infusion = 175 mL /hr

          Drop factor = 15 gtt /min

               gtt / min = 175 * 15 / 60

                                 = 43.75 gtt / min.

b. Infusion time

           Formula = Total volume (ml) / flow rate (ml/hr)

             Total volume = 2000 ml

             Infusion time     = 175 ml/hr

                               = 2000 / 175

                                = 11.43 hr.

c. Completion time

                 * It is the amount of time taken for completion of IV infusion.

                 * Infusion was started at 21:35.

                 * So the completion time = 21:35 pm + infusion time.

                                                        = 21.35 pm + 11.43 hrs

                                                           = 9.20 am